# New Exotic four-manifolds with $\mathbb{Z}/2\mathbb{Z}$ fundamental group

### v1

@@ Abstract.
We extend a construction of Stipsicz-Szabó ([13]) of infinitely many irreducible exotic smooth structures of some closed four-manifolds with even $b_{2}^{+}$ and fundamental group $\mathbb{Z}/2\mathbb{Z}$. We use the double node surgery and rational blow down constructions of Fintushel-Stern ([6, 4]) on some elliptic fibrations equipped with a free involution. The construction is done in an equivariant manner and the factor manifolds are distinguished by the Seiberg-Witten invariants of their universal covers.
@@ 1. Introduction
In this paper, we will expand the construction given in [13, Theorem 4.6] by exploiting the extra two $E(n)$ pieces of the fibration described below. The following theorem relies on the rational blow down construction, knot and double node surgery constructions of Fintushel-Stern (see [4, 5, 6] respectively).
@@ 2.2. Knot surgery
Consider a knot $K\subset S^{3}$, and a simply connected smooth manifold $X$ with $b_{2}^{+}>1$, and with an embedded homologically essential torus $T$ (in the following we denote the $n$-torus $T^{n}$) of self-intersection 00 and simply connected complement. The manifold $X_{K}:=(X\setminus\nu T)\cup_{\phi}(S^{3}\setminus\nu K)\times S^{1}$ is called a knot surgery of $X$ using $K$, where $\phi:T^{3}\to T^{3}$ is chosen so that the longitude of $K$ is identified with the normal circle of $T$. Note that this does not determine $\phi$ completely: we pick and fix such a function.
@@
In the special case of elliptic fibrations over the sphere, which we will be using, more can be done ([6]). Let $K$ be any genus one knot, and pick a minimal genus Seifert surface $\Sigma$ and a non-separating loop $\Gamma$ inside $\Sigma$ satisfying:
@@
With $K$ as previously and $T$ a smooth fiber in a double node neighborhood, we do knot surgery with the map $\phi$ picked so that the meridian of $K$ is identified with the vanishing cycle of the singular fibers (both are the same, since they have the same monodromy).
@@ 2.3. Rational blow down
The final ingredient is the rational blow down ([4, 10]). Denote by $C_{p}$ a linear plumbing (consecutive terms intersect transversely at a single point) of $p-1$ spheres with self-intersections $-(p+2),-2,\dots,-2$. The boundary of this linear plumbing is the lens space $\partial C_{p}=L(p^{2},p-1)$. The main observation is that there is a manifold $B_{p}$ with the same boundary, and with $H_{*}(B_{p};\mathbb{Q})=H_{*}(\{pt\};\mathbb{Q})$ ([7, Figure 8.42.]). If $C_{p}\subset X$ for some simply connected 4-manifold $X$, we call $X_{p}:=(X\setminus\nu C_{p})\cup B_{p}$ the rational blow down of $X$.
@@
$C_{p}$ is Seiberg-Witten tautly embedded if $|\alpha(u_{1})|\leq p$ (where $u_{1}\subset C_{p}$ is the sphere with self-intersection $-p-2$), and $\alpha(u_{i})=0$ (for the $-2$ spheres of the configuration) is satisfied for all basic classes $\alpha$.
@@
(1) $\frac{\alpha^{2}-(3\sigma(X)+2\chi(X))}{4}=0$
@@ 2.4. An involution on $E(2n+1)$
We briefly recall the construction of the extended involution, still denoted by $\iota:E(1)\to E(1)$ from [13, Section 3]. Consider $S^{2}\times S^{2}$ and the antipodal map on both factors, denoted by $p\times p$. Now take two fibers and two constant sections, which are not mapped into each other by the respective $p$’s, call the union of these 4 spheres $C_{0}$, and $C_{1}=(p\times p)(C_{0})$ is another set of 4 spheres.
@@
Note, that the involution $p\times p$ extends if we blow up $S^{2}\times S^{2}$ at pairs of these intersection points, thus getting a well defined fibration $S^{2}\times S^{2}\#8\overline{\operatorname{\mathbb{CP}}^{2}}\to S^{2}$, still equipped with an involution, which we still call $\iota$. This fibration is an elliptic fibration $E(1)$.
@@ 3. Construction
More precisely, consider the fibration $E(2n+1)$ described in Subsection 2.4. Choose a section $s\in\Gamma(E(2n+1))$, and consider the image $\iota(s)$ of that section under the involution map $\iota$ (they have self-intersection $-2n-1$). Choose two smooth fibers (again equivariantly) of our fibration, and do knot surgery along them using the twist knot $K_{m}$.
Figure 3. The elliptic surface $E(2n+1)$ seen as the fiber sum of $E(1)$ with two copies of $E(n)$.
Since $g_{3}(K_{m})=1$, the sections become genus two surfaces after the surgery, and since
(2) $1=(ab)^{6}=(a^{3}b)^{3}=a^{9}(b^{a^{6}}b^{a^{3}}b)$
in the monodromy group (see [14] for further details) we can assume that there are two $I_{4}$ fibers in our $E(1)$, positioned symmetrically with respect to $\iota$.
@@
According to [6] we can use these fibers to exchange the genera for two positive double points on both fibers and both sections. Blowing up the four double points, the sections now have self-intersection $-2n-9$.
Next, we do $2k$ additional double node surgeries along the trefoil knot (we choose the left-handed trefoil $K_{1}$) positioned symmetrically with respect to $\iota$ in the two $E(n)$ parts of our $E(2n+1)$ (note that a number of other knots would work, but we choose the trefoil to keep the computations of the Seiberg-Witten invariants simple). As before, $g_{3}(K_{1})=1$ and its Alexander polynomial is $\Delta_{K_{1}}=t-1+t^{-1}$, allowing us to apply additional double node surgeries to produce double points on the spheres and blow them up. The existence of the necessary $I_{4}$ fibers for this construction follows from the monodromy factorisation in the next paragraph. This lowers the self-intersection of both sections to $-2n-9-8k$.
Next we produce a singular $I_{8n-2}$ fiber, where the dual intersection graph of the fiber and the two sections looks like Figure 5.
@@
This can be achieved using a result of Korkmaz-Ozbagci ([9]). From this, we compute the following factorisation in the mapping class group of the twice punctured torus, where $\beta$ denotes a right handed Dehn twist along the longitude, and $\overline{\beta}$ is its inverse (see Figure 6):
@@
As we can see, this decomposition does not contain any $\alpha_{2}$. In the fibration corresponding to this decomposition, the two sections would intersect the same $(-2)$-curve of the singular fiber. In order to separate the two sections we also need the decomposition $\delta_{1}\delta_{2}=\alpha_{1}^{6}\alpha_{2}^{3}\beta^{\alpha_{1}^{4}\alpha_{2}^{2}}\beta^{\alpha_{1}^{2}\alpha_{2}}\beta$.
@@
Using the two relations, we get that
@@
Finally, we perform two rational blow downs, one on this configuration and one on its pair under $\iota$. The fibers were chosen to be longer than needed for a $C_{2n+7+8k}$; two additional spheres of the singular fiber obtained from this choice are used to make the configuration disjoint from its pair under $\iota$, and to guarantee that the resulting manifold is simply connected. The fundamental group of the complement is generated by a normal circle of one of the endpoints of the linear plumbing $C_{2n+7+8k}$, the longer chain guarantees a bounding disc to this generator in the complement inside $E(2n+1)$, and by Van Kampen’s theorem we get simple connectivity (the complement of a section and a fiber in $E(2n+1)$ is also simply connected [7, Section 3.1]).
Doing this procedure once requires an $I_{4}$ for the double node surgery, and the blowups force the configuration $C_{p}$ to be longer by $8$, so we use $a^{12}$ from the monodromy in total (from now on, $a$ is a right handed Dehn twist in the mapping class group of the torus of which $\alpha_{1}$ and $\alpha_{2}$ are lifts). By Equation (3), we can produce $a^{8n-2}$ in the monodromy factorization of $E(n)$. Therefore, we have $8n-2=2n+7+12k$, i.e. $k=\lfloor\frac{n}{2}-\frac{3}{4}\rfloor$ gives the maximum number of times we can perform the surgery. Note that this means the coefficient $8n-6k$ is equal to $65n+$ 6 if $n$ is even, and to $95n+$ 9 if $n$ is odd.
@@
This choice leaves some singular fibers to work with, by using the factorisation of Equation (2) for the double node surgeries, the results can be improved by about 1 extra knot surgery for every fifth value of $n$, but the formulas become much more convoluted.
@@
Topologically, $\chi(E(2n+1))=24n+12$, which we change by $4k+4-4n-12-16k$ with the blow ups and the rational blow downs to obtain $420n-12k+$ 4. Furthermore, $b_{2}^{+}(E(2n+1))=4n+1,\ b_{2}^{-}(E(2n+1))=20n+9$, we only add and remove negative definite submanifolds, so the signature becomes $4n+1-(20n+9+4k+4-4n-12-16k)=-12n+12k$. This number is not divisible by 16 by assumption, thus our manifolds are not spin i.e., type I. Factorisation by $\iota$ halves both $\sigma$ and $\chi$, so by the homeomorphism classification of $\pi_{1}=\mathbb{Z}/2\mathbb{Z}$ smooth manifolds the factor is homeomorphic to $Z_{1}\#2n\operatorname{\mathbb{CP}}^{2}\#(8n-6k)\overline{\operatorname{\mathbb{CP}}^{2}}$, but as detected by the Seiberg-Witten invariant of its universal cover, they are not diffeomorphic.
Finally by the above calculation and Equation (1), we see that $\alpha^{2}=\frac{3\sigma+2\chi}{4}=4n+12k+8$ for our basic class $\alpha$. Irreducibility follows from this, since in a reducible manifold there are two basic classes, the difference of which has square $-4$ by [12, Lemma 2.3], but in our case this number is always positive. ∎

### v2

@@ Abstract.
We extend a construction of Stipsicz-Szabó of infinitely many irreducible exotic smooth structures of some closed four-manifolds with even $b_{2}^{+}$ and fundamental group $\mathbb{Z}/2\mathbb{Z}$. We use the double node surgery and rational blow down constructions of Fintushel-Stern on some elliptic fibrations equipped with a free involution. The construction is done in an equivariant manner and the factor manifolds are distinguished by the Seiberg-Witten invariants of their universal covers.
@@ 1. Introduction
In this paper, we will expand the construction given in [13, Theorem 4.6] by decomposing the elliptic fibration
$E(2n+1)=E(n)\#_{f}E(1)\#_{f}E(n)$
and exploiting the $E(n)$ summands of the fibration. We use the rational blow down construction, knot and double node surgery constructions of Fintushel-Stern (see [4, 5, 6] respectively) in order to construct infinite families of exotic simply connected four manifolds. Those manifolds will come equipped with involutions, which will provide the exotica of the following theorem:
@@ 2.2. Knot surgery
Consider a knot $K\subset S^{3}$, and a simply connected smooth manifold $X$ with $b_{2}^{+}>1$, and with an embedded homologically essential torus $T$ of self-intersection 00 and simply connected complement. The manifold $X_{K}:=(X\setminus\nu T)\cup_{\phi}(S^{3}\setminus\nu K)\times S^{1}$ is called a knot surgery of $X$ using $K$, where $\phi:T^{3}\to T^{3}$ is chosen so that the longitude of $K$ is identified with the normal circle of $T$, and $T^{3}=S^{1}\times S^{1}\times S^{1}$ denotes the $3$-torus. Note that this does not determine $\phi$ completely: we pick and fix such a function.
@@
In the special case of elliptic fibrations over the sphere – which we will be using – more can be done ([6]). Let $K$ be any genus one knot, and pick a minimal genus Seifert surface $\Sigma$ and a non-separating loop $\Gamma$ inside $\Sigma$ satisfying:
@@
With $K$ as previously and $T$ a smooth fiber in a double node neighborhood, we do knot surgery with the map $\phi$ picked so that the meridian of $K$ is identified with the vanishing cycle of the nodal fibers (both are the same, since they have the same monodromy).
@@ 2.3. Rational blow down
Another ingredient is the rational blow down ([4, 10]). Denote by $C_{p}$ a linear plumbing (consecutive terms intersect transversely at a single point) of $p-1$ spheres with self-intersections $-(p+2),-2,\dots,-2$. The boundary of this linear plumbing is the lens space $\partial C_{p}=L(p^{2},p-1)$. The main observation is that there is a manifold $B_{p}$ with the same boundary, and with $H_{*}(B_{p};\mathbb{Q})=H_{*}(\{pt\};\mathbb{Q})$ ([7, Figure 8.42.]). If $C_{p}\subset X$ for some simply connected 4-manifold $X$, we call $X_{p}:=(X\setminus\nu C_{p})\cup B_{p}$ the rational blow down of $X$.
@@
$C_{p}$ is Seiberg-Witten tautly embedded if $|\alpha(u_{1})|\leq p$ (where $u_{1}\subset C_{p}$ is the sphere with self-intersection $-p-2$), and $\alpha(u_{i})=0$ (for the ($-2$)-spheres of the configuration) is satisfied for all basic classes $\alpha$.
@@
(1) $\frac{{\alpha^{2}([X])}-(3\sigma(X)+2\chi(X))}{4}=0$
@@
2.4. Monodromy factorisation.
Let $f:X\rightarrow Y$ be an elliptic fibration, and $\Delta=\{p_{1},\dots,p_{n}\}\subset Y$ be the discriminant locus of the fibration (i.e. the points whose fiber is singular). We assume that $Y$ is the sphere $S^{2}$. Let $p\in\Delta$, and restrict the fibration to a small circle around $p$, which is disjoint from $\Delta$. By going around the small circle, we get a self-diffeomorphism of the generic fiber $T^{2}$, defined up to isotopy and conjugation. In other words, we get an element defined up to conjugation in the mapping class group $\Gamma_{1}$ of $T^{2}$. We will call this conjugacy class the monodromy of the singular fiber $f^{-1}(p)$. The mapping class group of the torus admits a presentation:
$\Gamma_{1}=\langle a,b\ |\ aba=bab,(ab)^{6}=1\rangle,$
where $a$ and $b$ correspond to the right-handed Dehn twists around the natural generators of $H_{1}(T^{2})$. The monodromy of the singular fiber $I_{n}$ is conjugate to $a^{n}$ in this presentation. Throughout this paper we use the convention $a^{b}=\overline{b}ab$ where $a\overline{a}=1$.
Suppose now that we have a product of right-handed Dehn twists
$t_{1}\dots t_{m}=(t_{1}\dots t_{i_{1}})(t_{i_{1}+1}\dots t_{2})\dots(t_{i_{k}+1}\dots t_{m})\in\Gamma_{1}.$
This induces an elliptic fibration $X\rightarrow D^{2}$ having singular fibers $F_{j}$ ($j=1,\dots,k$), with monodromies conjugate to $(t_{i_{j}+1}\dots t_{i_{j+1}})$. If the product $t_{1}\dots t_{m}=1$ in $\Gamma_{1}$, then the induced fibration closes up to a fibration over $S^{2}$. Let $\Gamma_{1,k}$ denote the mapping class group of a compact, connected orientable genus $1$ surface with $k$ boundary components (the diffeomorphisms and isotopies are assumed to fix the boundary points). A lift of the relation $t_{1}\dots t_{m}=1$ to an identity in $\Gamma_{1,k}$ shows the existence of $k$ disjoint sections in the induced fibration. Let us denote by $t_{\delta_{i}}\in\Gamma_{1,k}$ the right-handed Dehn twist around a circle parallel to the $i$-th boundary component. If the lift to $\Gamma_{1,k}$ is of the form $t_{1}\dots t_{m}=t_{\delta_{1}}^{n_{1}}\dots t_{\delta_{k}}^{n_{k}}$ then the self-intersection of the $i$-th section is $-n_{i}$.
In our case we will look at products of the form $t_{1}\dots t_{12}=1$ in $\Gamma_{1}$, which will give us an elliptic fibration $E(1)\rightarrow S^{2}$. Taking the $n$-th power of the relation will give us an elliptic fibration $E(n)\rightarrow S^{2}$.
Remark 2.8.
From this description, it is clear that given a factorisation of right-handed Dehn twists $t_{1}^{n_{1}}\dots t_{m}^{n_{m}}=1$ one can use the identity $t^{n+m}=t^{n}t^{m}$ to obtain another elliptic fibration which has one more discriminant point. Geometrically, this means for example that one $I_{4}$ fiber can be switched out for two $I_{2}$ fibers, without changing the diffeomorphism type of the total space, and the structure of the fibration is only changed near the original singular fiber.
This shows that an $I_{n\times m}$ fiber can be replaced by $m$ number of $I_{n}$ fibers with the same monodromy. In particular, an $I_{4m}$ fiber can be replaced with $m$ number of $I_{4}$ fibers, and each $I_{4}$ can also be changed to $4$ nodal fibers, allowing two double node surgeries to be performed.
@@ 2.5. An involution on $E(2n+1)$
We briefly recall the construction of the extended involution, still denoted by $\iota:E(1)\to E(1)$ from [13, Section 3]. Consider $S^{2}\times S^{2}$ and the antipodal map on both factors, denoted by $p\times p$. Let $a,c\in S^{2}$ and $b,d\in S^{2}$ such that $c\neq p(a)$ and $d\neq p(b)$. Let $C_{0}=S^{2}\times\{a\}\cup S^{2}\times\{c\}\cup\{b\}\times S^{2}\cup\{d\}\times S^{2}$.
This gives us four spheres (two fibers and two sections) in $S^{2}\times S^{2}$. Now $C_{1}=(p\times p)(C_{0})$ is another configuration of four spheres with $|C_{0}\cap C_{1}|=8$.
$S^{2}$
$S^{2}$
$C_{0}$
$C_{1}$
$a$
$c$
$b$
$d$
@@
Note, that the involution $p\times p$ extends if we blow up $S^{2}\times S^{2}$ at pairs of these intersection points, thus getting a well defined fibration $S^{2}\times S^{2}\#8\overline{\operatorname{\mathbb{CP}}^{2}}\to S^{2}$, still equipped with an involution, which we still call $\iota$. This fibration is an elliptic fibration with total space $E(1)$.
@@ 3. Construction
More precisely, consider the fibration $E(2n+1)$ described in Subsection 2.5. Choose a section $s\in\Gamma(E(2n+1))$, and consider the image $\iota(s)$ of that section under the involution map $\iota$ (they have self-intersection $-(2n+1)$). Choose two smooth fibers (again equivariantly) of our fibration, and do knot surgery along them using the twist knot $K_{m}$.
$\mathbb{Z}/2\mathbb{Z}$-action
$\iota(s)$
$s$
Figure 3. Schematic picture of the elliptic surface $E(2n+1)$ seen as the fiber sum of $E(1)$ with two copies of $E(n)$.
Since $g_{3}(K_{m})=1$, the sections become genus two surfaces after the surgery, and following Section 2.4 we write
(2) $1=(ab)^{6}=(a^{3}b)^{3}=a^{4}a^{4}a(b^{a^{6}}b^{a^{3}}b)$
in the monodromy group (see [14, Section 2] for further details) we can assume that there are two $I_{4}$ fibers in our $E(1)$, positioned symmetrically with respect to $\iota$.
Remark 3.1.
We use the relation $aba=bab$ (c.f. Subsection 2.4) and then use conjugation $aba=a^{2}b^{a}$ several times to move the $a$’s to the required position.
$12m-$ 1
@@
Using Remark 2.8 and [6] we can perform double node surgeries on these $I_{4}$ fibers to exchange the genera for two positive double points on both sections. Indeed, with one $I_{4}$ fiber positioned near the knot surgery we can trade one genus for a positive double point on both sections. The same can be done with two $I_{4}$ fibers equivariantly: carrying out the first two double node surgeries will change one genus on both $s$ and $\iota(s)$ to a positive double point, and doing the same construction on their "mirror image" changes the remaining genus into a positive double point on both sections (see Remark 2.8). Note that the blow up of a positive double point decreases the self-intersection of the blown up section by $4$, and decreases the homology class of the section by $2E$, where $E$ is the exceptional sphere of the blow up. Blowing up the four double points, the sections now have self-intersection $-2n-9$.
Next, we do $4k$ additional double node surgeries along the trefoil knot (we choose the left-handed trefoil $K_{1}$) positioned symmetrically with respect to $\iota$ in the two $E(n)$ parts of our $E(2n+1)$.
Remark 3.2.
Note that a number of other knots would work, but we choose the trefoil to keep the computations of the Seiberg-Witten invariants simple.
As before, $g_{3}(K_{1})=1$ and its Alexander polynomial is $\Delta_{K_{1}}=t-1+t^{-1}$, allowing us to apply additional double node surgeries to produce double points on the sections and blow them up.
This lowers the self-intersection of both sections to $-2n-9-8k$. Thus we need to find $k$ many $I_{4}$ fibers in $E(n)$ for this construction. These fibers and their images under $\iota$ will be used to do the double node surgeries mentioned. We produce a singular $I_{8n+1}$ fiber, where the dual graph of the fiber and the two sections looks like Figure 5.
$s$
$\iota(s)$
$38n-$ 3
@@
The first equality of the next equation is a result of Korkmaz-Ozbagci ([9, Section 3.2]). From this, we compute the following factorisation in the mapping class group of the torus with two discs removed, where $\beta$ denotes a right handed Dehn twist along the longitude, and $\overline{\beta}$ is its inverse (see Figure 6), using the same technique mentioned in Remark 3.1 to obtain the last two equalities:
@@
Remark 3.3.
Note that Dehn twists along disjoint curves commute, hence $\alpha_{1}$ and $\alpha_{2}$ commute, and we can use cyclic permutation since $\alpha_{1}$, $\alpha_{2}$ and $\beta$ commute with $\delta_{1}$ and $\delta_{2}$.
As we can see, this decomposition does not contain any $\alpha_{2}$ (without conjugation). This means that in the fibration corresponding to this decomposition, the two sections would intersect the same $(-2)$-curve of the singular fiber. In order to separate the two sections we also need the decomposition $\delta_{1}\delta_{2}=\alpha_{1}^{6}\alpha_{2}^{3}\beta^{\alpha_{1}^{4}\alpha_{2}^{2}}\beta^{\alpha_{1}^{2}\alpha_{2}}\beta$ which follows from the same principles as the previous.
$\alpha_{1}$
$\alpha_{2}$
$\delta_{1}$
$\delta_{2}$
$\beta$
$\delta_{1}$
$\delta_{2}$
$28n-$ 2
$\alpha_{1}$
$\alpha_{2}$
@@
Multiplying $n-1$ times the first decomposition with the second and using the method outlined in Section 2.4, we obtain an identity of the form
@@
Consider the union of $s$, and the ($-2$)-curves of the just constructed $I_{8n+1}$ and its pair under $\iota$ (see Figure 7). Since $s$ has self-intersection $-(2n+8k+9)$, this contains a $C_{2n+7+8k}$ subconfiguration, which is disjoint from its pair under $\iota$ since we have two additional spheres in the fiber. The fundamental group of the complement is generated by a normal circle of one of the endpoints of the linear plumbing $C_{2n+7+8k}$, the longer chain guarantees a bounding disc to this generator in the complement inside $E(2n+1)$. Since $E(2n+1)$ is simply connected, any curve in $E(2n+1)\setminus(C_{2n+7+8k}\cup\iota(C_{2n+7+8k}))$ bounds a disc in the ambient space. This intersects $\partial C_{2n+7+8k}=L((2n+7+8k)^{2},2n+6+8k)$ in some arcs. In general, lens spaces have abelian fundamental group, so it is enough to calculate homologically. These intersection arcs can be homotoped to the end of the chain, where they bound a disc.
$s$
$\iota(s)$
$\mathbb{Z}/2\mathbb{Z}$-action
$38n-$ 3
$38n-$ 3
Figure 7. The discs in the complement of the configuration, and the punctured $S^{2}$’s (represented by thickened lines).
Consider one of the $E(n)$ submanifolds and its pair under $\iota$. We wish to compute a bound on $k$ and to this end give an overview of our construction. It is enough to make the computation in one of the $E(n)$ parts of $E(2n+1)$, since its pair under $\iota$ is disjoint from it, we can easily extend everything equivariantly. To increase $k$ by one, we need an $I_{4}$ fiber to do two double node surgeries, and after blowing up each section once, their framing drops by $8$, forcing us to construct a $C_{p+8}$ instead of a $C_{p}$ (the $C_{p}$ where the first vertex is $s$ has its $-2$ curves in $E(n)$, and its pair has its $-2$ curves in $\iota(E(n))$). Thus all together increasing $k$ once requires $\alpha_{1}^{12}$ and when $k=0$ we need an $\alpha_{1}^{2n+7}$, since the double node surgeries were done in the "middle" $E(1)$ part, where the involution acts non-trivially. By Equation (3), we can produce $\alpha_{1}^{8n-2}$ in the monodromy factorization of $E(n)$. Therefore, we have
(4) $8n-2=2n+7+8k+4k,$
i.e. $k=\lfloor\frac{n}{2}-\frac{3}{4}\rfloor$ gives the maximum number of times we can perform the surgery. Note that this means the coefficient $8n-6k$ is equal to $65n+$ 6 if $n$ is even, and to $95n+$ 9 if $n$ is odd.
@@
In the process, each double node surgery needs a pair of Lefschetz thimbles (corresponding to parallel vanishing cycles), and different surgeries need separate thimbles to achieve that the end-result (after the blow-ups) is embedded, rather than immersed. This explains the need of the large number of parallel vanishing cycles in our construction, and in turn the coefficient of $k$ in Equation (4).
@@
Topologically, $\chi(E(2n+1))=24n+12$, which we change by $4k+4-4n-12-16k$ with the blow ups and the rational blow downs to obtain $420n-12k+$ 4. Furthermore, $b_{2}^{+}(E(2n+1))=4n+1,\ b_{2}^{-}(E(2n+1))=20n+9$, we only add and remove negative definite submanifolds, so the signature becomes $4n+1-(20n+9+4k+4-4n-12-16k)=-12n+12k$. This number is not divisible by 16 by assumption, thus our manifolds are not spin i.e., type I. Taking the quotient by $\iota$ halves both $\sigma$ and $\chi$, so by the homeomorphism classification of $\pi_{1}=\mathbb{Z}/2\mathbb{Z}$ smooth manifolds the factor is homeomorphic to $Z_{1}\#2n\operatorname{\mathbb{CP}}^{2}\#(8n-6k)\overline{\operatorname{\mathbb{CP}}^{2}}$, but as detected by the Seiberg-Witten invariant of its universal cover, they are not diffeomorphic.
Finally by the above calculation and Equation (1), we see that $\alpha^{2}={3\sigma+2\chi}=4n+12k+8$ for our basic class $\alpha$. Irreducibility follows from this, since in a reducible manifold there are two basic classes, the difference of which has square $-4$ by [12, Lemma 2.3], but in our case this number is always positive. This shows that the universal cover of our manifolds are irreducible, and thus the quotient is also irreducible. ∎